OpenStudy (destinykiara99):

Find a cubic function with the given zeros. Square root of six., - Square root of six., -3

OpenStudy (destinykiara99):

@Mehek14 @ganeshie8 @mayankdevnani

sam (.sam.):

If \(\sqrt{6}\), \(\sqrt{6}\), \(3\) are the zeros of a cubic function, then \(a=\sqrt{6}\), \(b=\sqrt{6}\), \(c=3\) \[(x-a) ( x-b) ( x - c ) = 0\] \[(x-\sqrt{6}) ( x-\sqrt{6}) ( x - 3 ) = 0\]

sam (.sam.):

Erm \(b=-\sqrt{6}\) sorry

OpenStudy (destinykiara99):

It's \[\sqrt{6}, -\sqrt{6}, -3 \] not both positive

sam (.sam.):

And the -3 as well

sam (.sam.):

Yeah

OpenStudy (destinykiara99):

Lol yeah :)

sam (.sam.):

Just use this and expand it and you'll get your cubic function.

OpenStudy (destinykiara99):

uhm okayy

sam (.sam.):

Let's see \[(x−\sqrt{6})(x+\sqrt{6})(x+3)=0 \\ \\ \]

sam (.sam.):

\[(x−\sqrt{6})(x+\sqrt{6})(x+3)=0 \\ \\ (x^2-6)(x+3)=0 \\ \\ x^3+3x^2-6x-18=0\]

OpenStudy (destinykiara99):

Oh! I get it now! Thanks!!

sam (.sam.):

Yw :)